# Measure Twice, Match Once: Flexible Graph Similarity Computation With A Proactive Optimization Strategy

### v1

@@ 5.1.5. Evaluation Metrics
Table 2. Effectiveness of models on test sets. We bold the best and underline the second models. ’OOM’ indicates out of memory.
@@
5.2. Results
5.2.1. Effectiveness
In this experiment, we compared the performance of our proposed Gen against all baselines under a uniform cost setting (operation costs set to 1). Results for the VJ and Beam algorithms on IMDBMulti were excluded, as IMDBMulti’s training targets are the smallest distances calculated by these algorithms. As shown in Table 2, Gen consistently ranks first or second across all metrics on the three datasets, demonstrating its competitive performance. GNN-based methods generally outperform conventional ones, as the former leverage enriched information extracted by GNNs. Graph comparison methods, such as Greed and Eric, outperform most node comparison competitors. This suggests that there is no clear winner between the two categories in this setting, where all models predict inherent graph differences. Most node comparison methods underperformed due to their reliance on node-level guidance, which fails to evaluate matches from a graph perspective. Among these methods, GraphEdX demonstrated advantages on the node-labeled AIDS700nef dataset by using XOR operation as a node distance measure. However, it underperformed on the LINUX dataset because it is unlabeled. Additionally, it encountered out-of-memory errors on IMDBMulti with larger graphs because it, by design, requires padding all graphs to the same size, whereas other models, such as GEDGNN and Gen, pad the graphs within graph pairs instead. GMN-match iteratively fuses and propagates node representations across graphs. It showed competitive performance on two datasets but suffered from error accumulation due to its greedy node-level fusion. Similarly, GraphSim’s BFS-based matrix reordering led to instability, as indicated by large standard deviations. In contrast to these node comparison methods, Gen introduces a proactive optimization strategy that leverages node alignment difficulty as guidance and optimizes it by considering interdependencies between matches. This strategy enables Gen to achieve up to a 37.8% error reduction in RMSE and a 30% reduction in MAE on the LINUX dataset. Additionally, Gen generally improves ranking performance by up to 4.1%, with a marginal decrease of up to 0.4% in Kendall’s tau $\tau$ on the IMDBMulti dataset. This suggests potential optimization for datasets like IMDBMulti, whose graphs have a low diameter of up to 2. A case study comparing the ranking results of Gen against the ground truth provided by A* is available in Figure 3.
Figure 3. Case Study of ranking results on AIDS700nef (labeled) and IMDBMulti (unlabeled).
@@ 5.2.1. Effectiveness
Table 3. We add ’Rank’ to indicate the current ranking based on RMSE and show the improvements over the ranking for the AIDS700nef dataset in Table 2 under a uniform cost setting, as shown in parentheses.
@@
Figure 4. Time Efficiency of GSC models. The bars with black edges represent node comparison-based methods.
5.2.2. Time Efficiency
In this experiment, we evaluated the training and inference times of models. Results for VJ and Beam were omitted, as prior experiments showed learning-based methods outperform them in efficiency (Bai et al., 2019; Zhang et al., 2021). Figures 4(a) and 4(b) illustrate the average time and epochs needed for convergence. Our proposed Gen ranks fourth or fifth fastest in training time, outperforming all node comparison (NC) models, and competing with the slowest graph comparison (GC) model, Eric, which constrains node-graph alignment during training. Although SimGNN shares a similar batch runtime with ours, it requires more epochs to converge. Figure 4(c) illustrates the total inference time on testset, GC methods are generally more time-efficient than NC ones, all of latter exhibit at least quadratic time complexity. Compared with the fastest NC method, SimGNN, Gen incurs an 11.2% increase in inference time but achieves up to 80.6% error reduction. GEDGNN and GraphEdX are among the slowest, as they establish mappings in a matrix formulation based on matching costs and require 20 refinement iterations to ensure performance. Similarly, GMN-match requires up to 5 iterative refinements for comparable accuracy. These refinements significantly increase computational demand as the number and size of graph pairs grow, highlighting by their increased time consumption on the LINUX and IMDBMulti datasets. Compared with the most relevant and competitive GMN-match, Gen reduces inference time by up to 72.7%. This improvement comes from the effective graph perspective guidance provided our proposed strategy, which enable Gen to make more informed decisions, minimizing the need for refinements, and leading to faster training and inference.
5.2.3. Adaptation to Varying Costs
In this experiment, we assess the adaptability of models to non-uniform and varying cost settings using the node-labeled AIDS700nef dataset. We evaluate two cost settings: Setting 1 assigns costs of 1 for node substitution, 2 for node deletion, 1 for node insertion, 3 for edge deletion, and 1 for edge insertion, denoted as [1, 2, 1, 3, 1]. In Setting 2, the corresponding costs are [3, 2, 3, 0, 2]. The results, presented in Table 3, show that: (i) Under non-uniform cost settings: Our proposed Gen model demonstrates superior adaptability, achieving the best performance across all metrics for both settings. The performance of Greed degraded sharply, likely due to its use of symmetric Euclidean distance, whereas under these settings, the matching cost from $s$ to $t$ may differ from that of $t$ to $s$, making it asymmetric. GMN-match also underperformed, as it greedily fuses nodes based on inherent node similarity, which can lead to significant error accumulations. Notably, all ’Rank’ improvements occurred for node comparison methods, suggesting that they are more suitable for flexible GED computation. (ii) Under varying costs: We trained Gen mix, TaGSim mix, and GraphEdX mix, these consider operation types and costs, using a mixed dataset under two cost settings. Results show that Gen mix substantially outperforms all models trained under a single cost setting and TaGSim mix and GraphEdX mix trained under varying settings, demonstrating better adaptability. The sharp performance degradation of TaGSim mix and GraphEdX mix highlights their key limitation, as they treat operation costs as irrelevant when determining the mappings, resulting in static mappings that are suboptimal under non-uniform settings. In contrast, the improvement of Gen mix over Gen demonstrates that our principled solution enhances the model’s understanding of GED computation. By preserving consistent node distances and dynamically adjusting them based on specific operation costs, it enables better adaptation to varying cost settings.
Table 4. Ablations Study on our proposed components.
Dataset AIDS700nef LINUX IMDBMulti
Gen wo interdependencies $1.101_{\pm\text{0.041}}$ $0.905_{\pm\text{0.006}}$ $0.296_{\pm\text{0.034}}$ $\underline{0.990_{\pm\text{0.005}}}$ $14.391_{\pm\text{2.568}}$ $0.982_{\pm\text{0.004}}$
5.2.4. Ablation Study
In this experiment, we evaluate the effectiveness of our proposed strategy. We test three variants: (1) Gen wo global: This variant replaces the global cost ratio with absolute node matching costs to estimate alignment difficulties. (2) Gen wo interdependencies: This variant reweighs matches solely based on their global cost ratios, where higher ratios result in lower weights. (3) Gen wo intra-graph: This variant does not compute the neighborhood alignment difficulty and considers inter-graph interdependency based only on node alignment difficulties. Note that the inter-graph interdependency cannot be ablated separately as it produces the decision for each match. As shown in Table 4, Gen wo interdependencies greedily reweighs matches based on global cost ratios, leading to a significant performance drop across all datasets, with RMSE increasing by up to 91.1%. On the other hand, Gen wo global incurs a smaller increase, at most 25%, indicating that considering interdependencies is more crucial than the global cost ratio. Meanwhile, ignoring the difficulties of neighboring nodes has a smallest impact. Gen wo intra-graph showcased comparable performance on AIDS700nef and LINUX, yet incurred a 12.1% increase in RMSE on IMDBMulti. This is likely because the matches in IMDBMulti are more ambiguous, making the neighborhood alignment difficulty a crucial factor.
Figure 5. Hyperparameter sensitivity.
5.2.5. Hyperparameter Sensitivity

### v2

@@ 5.4 Complexity Analysis
Table 1: Effectiveness of models on test sets. We bold the best and underline the second models. ‘OOM’ indicates out of memory.
@@
6 Evaluations
In this section, we evaluate Gen against twelve baselines, including two conventional inexact algorithms, on three commonly used GED datasets and synthetic graphs, focusing on Effectiveness, Time Efficiency, Adaptation to Varying Cost Settings, and Performance on Larger Synthetic Graphs. The hyperparameter senstitivity analysis are available in Appendix D.4.
6.1 Experimental Setup
Dataset. We conduct experiments on three commonly used GED datasets: AIDS700nef, Linux and IMDBMulti. Following prior work [16, 3, 20], we generate synthetic graphs with a node size ranging from 100 to 400, as extreme data to to assess model performance on larger graphs. We generate 51,200 graphs for training and another 38,400 graphs in three node size groups: $[100,200]$, $[200,300]$, and $[300,400]$, for testing. The detailed dataset descriptions can be found in Appendix C.1.
Baselines. We include representative, recent, and state-of-the-art GED computation models and approximation algorithms as baselines: (1) Graph Comparison Methods: GMN-emb [16], TaGSim [3], Greed [22] and Eric [35]; (2) Node Comparison Method: GMN-match [16], SimGNN [4], GraphSim [5], GOTSim [9], GEDGNN [20] and GraphEdX [13]. (3) Approximate Algorithms: Volgenant–Jonker (VJ) algorithm [10] and Beam [19]. We generally use the official implementations and hyperparameters provided by the authors. More details about the baselines and the implementation of Gen can be found in Appendix C.2 and C.3.
Experimental Protocol. We partitioned the datasets into training and test graphs at a 4:1 ratio, with 20% of the training graphs used for validation. Training pairs are sampled from the training set. All models are trained for ten warm-up epochs and evaluated on the validation set every epoch. Early stopping with a patience of 50 is applied to prevent overfitting and ensure that each model is sufficiently trained. More details are available in Appendix C.4.
Evaluation Metrics. We evaluate models’ prediction error and ranking ability. For error metrics, we report Root Mean Square Error (RMSE) and Mean Absolute Error (MAE). For ranking, we use Spearman’s Rank Correlation Coefficient ($\rho$), Kendall’s Rank Correlation Coefficient ($\tau$), and Precision at $k$ ($p@k$). Details of metric definitions are provided in Appendix C.5. We report the average and standard deviation of results across five runs.
6.2 Results
Effectiveness. We compare the performance of Gen against all baselines under a uniform cost setting (operation costs set to 1). Results for VJ and Beam on IMDBMulti are excluded, as the training targets of this dataset are the smallest distances produced by these algorithms. As shown in Table 1, Gen consistently ranks first or second across all metrics and datasets, demonstrating strong overall performance. GNN-based methods generally outperform conventional ones due to their ability to encode richer structural information. Notably, graph comparison methods (e.g., Greed, Eric) outperform most node comparison approaches, suggesting no definitive advantage between the two paradigms under this setting. Gen clearly outperforms most node comparison methods, as they often rely on isolated node-level distances, while Gen proactively optimizes guidance to capture graph-level matching dependencies, enabling informed one-shot matching. Consequently, Gen achieves up to a 37.8% reduction in RMSE and a 30% reduction in MAE on LINUX, and improves ranking performance by up to 4.1%, with only marginal declines ($\leq$0.4%) in a few cases. Further discussions and and a ranking case study can be found in Appendix D.1 and D.2, respectively.
@@ 6.2 Results
Table 2: ‘Rank’ indicates the current ranking based on RMSE and show the improvements over the ranking for the AIDS700nef dataset in Table 1 under a uniform cost setting, as shown in parentheses.
@@
Time Efficiency. We evaluate the training and inference time of all models. Results for VJ and Beam are excluded, as prior studies have shown that learning-based methods significantly outperform them in efficiency [4, 31]. As shown in Figure 5, our proposed Gen ranks fourth or fifth fastest in training time, outperforming all node comparison (NC) models and closely approaching the slowest graph comparison (GC) baseline, Eric. While SimGNN shows similar batch runtimes, it requires more training epochs to converge. At inference, GC methods are generally more time-efficient than NC ones, the latter exhibiting at least quadratic complexity. Compared with the most relevant and competitive NC baseline, GMN-match, Gen reduces inference time by up to 72.7% while achieving comparable or better performance. This efficiency gain stems from our proactive optimization strategy, which captures graph-level dependencies and reduces the need for iterative refinements.
Adaptation to Varying Costs. We assess the adaptability of models to non-uniform and varying cost settings using the node-labeled AIDS700nef dataset. We evaluate two cost settings: Setting 1 assigns costs of 1 for node substitution, 2 for node deletion, 1 for node insertion, 3 for edge deletion, and 1 for edge insertion, denoted as [1, 2, 1, 3, 1]. In Setting 2, the corresponding costs are [3, 2, 3, 0, 2]. As shown in Table 2, (i) under non-uniform cost settings: Gen achieves the best performance in both settings, showing strong adaptability to asymmetric edit costs. In contrast, Greed performs poorly due to its use of symmetric distance measure, and GMN-match suffers from error accumulation caused by its greedy local fusion mechanism. Notably, rank improvements are more prominent among NC models compared with GC ones, indicating NC models’ suitability for flexible GED estimation. (ii) Under varying costs: We further train Gen mix, TaGSim mix, and GraphEdX mix on a dataset combining both settings. Gen mix substantially outperforms all baselines, demonstrating better adaptability. The sharp performance degradation of TaGSim mix and GraphEdX mix highlights their key limitation, as they treat operation costs as irrelevant during aligning stage, they produce static mappings that are suboptimal under non-uniform settings. In contrast, the improvement of Gen mix over Gen highlights the robustness and generalization capability of our principled solution. This indicates that exposure to diverse cost settings does not impair Gen’s performance, but instead facilitates a deeper understanding of GED computation, leading to more accurate predictions.
Table 3: Ablations Study on our proposed components.
Dataset AIDS700nef LINUX IMDBMulti
Gen wo dependencies $1.101_{\pm\text{0.041}}$ $0.905_{\pm\text{0.006}}$ $0.296_{\pm\text{0.034}}$ $\underline{0.990_{\pm\text{0.005}}}$ $14.391_{\pm\text{2.568}}$ $0.982_{\pm\text{0.004}}$
Ablation Study. We evaluate the effectiveness of our proposed strategy. We test three variants: (1) Gen wo global: This variant replaces the global cost ratio with absolute node matching costs to estimate alignment difficulties. (2) Gen wo dependencies: This variant reweighs matches solely based on their global cost ratios, where higher ratios result in lower weights. (3) Gen wo intra-graph: This variant does not consider intra-graph connections. Note that the inter-graph discrepancies cannot be ablated separately as they produce the matching decisions. As shown in Table 3, Gen wo dependencies greedily reweighs matches based on global cost ratios, leading to a significant performance drop across all datasets, with RMSE increasing by up to 91.1%. On the other hand, Gen wo global incurs a smaller increase, at most 25%, indicating that considering dependencies is more crucial than the global cost ratio. Meanwhile, ignoring the difficulties of neighboring nodes has a smaller impact. Gen wo intra-graph showcased comparable performance on AIDS700nef and LINUX, yet incurred a 12.1% increase in RMSE on IMDBMulti. This is likely because the matches in IMDBMulti are more ambiguous, making the neighborhood alignment difficulty a crucial factor.
@@
Figure 4: Time Efficiency of GSC models. The bars with black edges represent node comparison-based methods.
Figure 5: Performance of models on larger synthetic graphs across diverse node size groups.
Performance on Larger Synthetic Graphs. GED is typically applied to small graphs, as its exact computation becomes challenging for graphs with more than 16 nodes [6]. To evaluate the performance of GSC models under extreme conditions, we followed [20] and assessed the performance of GSC models on larger synthetic graphs. We compared GED predictions among baseline methods and our proposed Gen against ground-truth GEDs. Figure 5 illustrates the results for the best NC and GC models during this experiment, GEDGNN and TaGSim, against our proposed Gen. The results show that Gen achieved the most accurate predictions, with data points closely clustered around the line representing perfect alignment between predictions and GEDs across all node size groups. The full results can be found in Appendix D.3.
7 Conclusion
This paper presents Graph Edit Networks (Gen), a learning-based model that captures the inherent flexibility of graph edit distance (GED) computation with an efficient proactive optimization strategy. We highlight a key limitation in existing methods in handling flexible GED computation: they neglect or treat operation costs as irrelevant when determining the mappings, resulting in static mappings across cost settings. To address this issue, we follow the GED computation principle and propose incorporating operation costs before establishing mappings. Instead of reactively refining mappings with insufficient node-level distances, we introduce a novel proactive optimization strategy. This strategy enables effective guidance capturing graph-level matching dependencies, allowing accurate one-shot match selections. Experiments on both real-world and synthetic datasets validate the model’s effectiveness, efficiency, and adaptability. Our proposed approach can be applied in graph retrieval scenarios, where Gen can return the top $k$ relevant graphs based on a query or rerank candidate graphs generated by fast graph comparison models to improve precision.
Appendix A Comparison Table
To further highlight the key differences in guidance and refinement strategies, we summarize the comparisons in Table 4.
Appendix B Pseudo Code for Gen
The pseudo code for Gen can be found in Algorithm 1.
Algorithm 1 The Gen Architecture
0:  A padded graph pair $(S,T)$ each has N nodes, where $S=(\mathbf{X}_{S},\mathcal{A}_{S})$, query graph $T=(\mathbf{X}_{T},\mathcal{A}_{T})$, the customized operation cost matrix $\mathcal{C}\in\mathcal{R}^{N^{2}\times|op|}$
0:  The estimated GED value between $S$ and $T$
1:  $\mathbf{X}^{0}_{S}=\operatorname{Linear}_{pre}(\mathbf{X}_{S})\in\mathbb{R}^{|\mathcal{V}|_{S}\times d_{hidden}}$ Initialize Node Embedding
2:  $\mathbf{X}^{0}_{T}=\operatorname{Linear}_{pre}(\mathbf{X}_{T})\in\mathbb{R}^{|\mathcal{V}|_{T}\times d_{hidden}}$
3:  for $l=1,\dots,K$ do
4:     $\mathbf{X}_{S}^{l}=\operatorname{GNN}(\mathbf{x}_{s}^{l-1}\in\mathbf{X}_{S}^{l-1},\mathcal{A}_{S})\in\mathbb{R}^{|\mathcal{V}_{S}|\times d_{hidden}}$ GNN Encoder
5:     $\mathbf{X}_{T}^{l}=\operatorname{GNN}(\mathbf{x}_{t}^{l-1}\in\mathbf{X}_{T}^{l-1},\mathcal{A}_{T})\in\mathbb{R}^{|\mathcal{V}_{T}|\times d_{hidden}}$
6:  end for
7:  $\mathbf{X}_{S}=\operatorname{MLP}(\mathbf{X}_{S}^{k})\in\mathbb{R}^{|\mathcal{V}_{S}|\times d}$ Reduce Dimensionality
8:  $\mathbf{X}_{T}=\operatorname{MLP}(\mathbf{X}_{T}^{k})\in\mathbb{R}^{|\mathcal{V}_{T}|\times d}$
9:  $\mathbf{H}_{S}=\text{Concat}(\mathbf{X}_{S},\mathbf{X}_{S}^{k})\in\mathbb{R}^{|\mathcal{V}_{S}|\times(|\mathcal{X}|+d)}$
10:  $\mathbf{H}_{T}=\text{Concat}(\mathbf{X}_{T},\mathbf{X}_{T}^{k})\in\mathbb{R}^{|\mathcal{V}_{T}|\times(|\mathcal{X}|+d)}$
11:  $\text{dist}(S,T)=\mathbf{H}_{S}-\mathbf{H}_{T}\in\mathbb{R}^{N^{2}\times(|\mathcal{X}|+d)}$ Cost-aware Scoring
12:  $c(S,T)=\text{Estimator}(\text{Concat}(\mathcal{C},\text{dist}(S,T)))\in\mathbb{R}^{N^{2}\times d}$
13:  for $c(s,t)\in c(S,T)$ do
14:     $\gamma(s,t)=exp(c(s,t))/c(S,T)\in\mathbb{R}^{d}$ Global Cost Ratios Initialization
15:  end for
16:  for $s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}$ do
17:     $\tilde{\gamma}(s)=\sum_{i\in\mathcal{V}_{T}}\gamma(s,i)\in\mathbb{R}^{d}$ Node Alignment Difficulty Computation
18:     $\tilde{\gamma}(t)=\sum_{i\in\mathcal{V}_{S}}\gamma(i,t)\in\mathbb{R}^{d}$
19:  end for
20:  for $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T}$do
21:     $\hat{\gamma}(v)=\text{MLP}\bigg{(}\text{Sum}(\tilde{\gamma}(v),\text{Sum}(\tilde{\gamma}(v^{\prime}):\mathcal{A}(v,v^{\prime})=1))\bigg{)}\in\mathbb{R}^{d}$ Neighborhood Alignment Difficulty Computation
22:  end for
23:  $\gamma(S)=\text{MLP}\bigg{(}\text{Concat}(\hat{\gamma}(S),\mathbf{H}_{S})\bigg{)}\in\mathbb{R}^{|\mathcal{V}_{S}|\times d}$ Add Residual Info
24:  $\gamma(T)=\text{MLP}\bigg{(}\text{Concat}(\hat{\gamma}(T),\mathbf{H}_{T})\bigg{)}\in\mathbb{R}^{|\mathcal{V}_{T}|\times d}$
25:  $\xi(S,T)=\text{MLP}\bigg{(}\text{Concat}(\gamma(S),\gamma(T))\bigg{)}\in\mathbb{R}^{N^{2}\times d}$ Compute Certainty
26:  for $s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}$ do
27:     $c(s)=\sum_{i\in\mathcal{V}_{T}}\xi(s,i)\cdot c(s,i)\in\mathbb{R}^{d}$ Compute Certainty-weighted Cost
28:     $c(t)=\sum_{i\in\mathcal{V}_{s}}\xi(i,t)\cdot c(i,t)\in\mathbb{R}^{d}$
29:  end for
30:  $c(S)=\sum_{s\in\mathcal{V}_{S}}c(s)\in\mathbb{R}^{d}$ Compute The Total Matching Costs
31:  $c(T)=\sum_{t\in\mathcal{V}_{T}}c(t)\in\mathbb{R}^{d}$
32:  $\text{GED}(S,T)=\text{Linear}\bigg{(}\text{Mean}(c(S),c(T))\bigg{)}\in\mathbb{R}^{1}$
33:  return $\text{GED}(S,T)$ Predict GED
B.1 GNN Encoder
$\mathbf{x}^{l}_{v}=\psi\bigg{(}\mathbf{x}^{l-1}_{v},\phi(\mathbf{x}^{l-1}_{u}:\mathcal{A}(v,u)=1)\bigg{)}$ (15)
Where $\phi$ is an aggregation function, and $\psi$ is a combination function, typically permutation-invariant operations like Sum, Mean, or Max. $\mathbf{x}^{l}_{v}$ and $\mathbf{x}^{l-1}_{v}\in\mathbb{R}^{d_{hidden}}$ represent the output of node $v$ from layers $l$ and $l-1$, respectively. $\mathbf{x}^{l-1}_{u}\in\mathbb{R}^{d_{hidden}}$ represents the output of $v$’s neighbor at the previous layer. The final layer representation $\mathbf{x}^{k}_{v}$ is then passed through a Multilayer Perceptron (MLP) to reduces the dimensionality of the representation to $d$. By altering GNNs with the ones considering edge information, this method can be extended to edge-labeled graphs without losing generalization.
Appendix C Experimental Setup
C.1 Dataset
We conduct experiments over three commonly used GED datasets:
  • AIDS700nef: 700 node labeled chemical compound structures with ten or less than ten nodes;
  • Linux: 1,000 unlabeled Program Dependence Graphs (PDG) generated from the Linux kernel, with ten or less than ten nodes;
  • IMDBMulti: 1500 unlabeled ego-networks of movie actors/actresses, with up to 89 nodes.
The graphs within the datasets and the exact or approximate (for IMDBMulti dataset) GED values between all possible graph pairs under a uniform cost setting are collected and provided by [4], resulting in abundant graph pairs. With the same graphs in the AIDS700nef dataset, we compute the GEDs under two non-uniform settings with the A*-based algorithm [2]. Note that exact GED computation is already challenging for graphs with 16 nodes [6]. Additionally, we generate synthetic graphs, among Erdős-Rényi, power-law, and Watts–Strogatz small-world graphs with a node size ranging from 100 to 400, as extreme data to further evaluate the effectiveness of models under the impact of graph size. For each synthetic graph, we apply $\Delta$ distinct operations to generate a target graph. The number of operations ranges from 10-30 percent of edge numbers. Following previous work [16, 3, 20], we regard this $\Delta$ as the ground-truth GED for synthetic graph pairs. We generate 51,200 graphs as training sets and another 38,400 graphs within three node size groups, including $[100,200]$, $[200,300]$, and $[300,400]$, as test sets. The detailed statistics of graphs and distributions of GEDs of these datasets can be found in Table 5 and Figure 6, respectively.
Figure 6: The GED distributions of datasets.
Table 5: Statistics of datasets.
Dataset #Graphs #Pairs #Testing Pairs #Labels Avg. #nodes Avg. #edges Max. #nodes Max. #edges Min. #nodes Min. #edges
AIDS700nef 700 490K 78,4K 29 8.90 8.80 10.00 14.00 2.00 1.00
LINUX 1000 1M 160k 1 7.58 6.94 10.00 13.00 4.00 3.00
IMDBMulti 1500 2.25M 360k 1 13.00 65.94 89.00 1467.00 7.00 12.00
Synthetic 51,200 25,6 K - 1 230.24 1789.02 400.00 8552.00 100.00 100.00
[100, 200] 12,800 6.4K - 1 148.61 1328.51 200.00 5806.00 100.00 196.00
[200, 300] 12,800 6.4K - 1 247.45 1979.85 300.00 7638.00 200.00 396.00
[300, 400] 12,800 6.4K - 1 348.14 2295.15 400.00 8083.00 300.00 100.00
C.2 Baselines
We include representative, recent, and state-of-the-art GED computation models as baselines: (1) Graph Comparison Methods: GMN-emb [16], TaGSim [3], Greed [22] and Eric [35]; (2) Node Comparison Method: GMN-match [16], SimGNN [4], GraphSim [5], GOTSim [9], GEDGNN [20], GraphEdX [13]. Additionally, we compare with well-known approximate algorithms, such as the Volgenant–Jonker (VJ) algorithm [10] and Beam [19]. Hungarian [23] is not included as it has been incorporated in GOTSim. Specifically:
  • VJ [10]: Volgenant–Jonker (VJ) algorithm is a linear sum assignment that greedily performs bipartite graph matching to minimize costs between nodes of two graphs.
  • Beam [19]: Beam search employs a best-first strategy, expanding partial solutions with the lowest-cost matches at each step to approximate GED.
  • SimGNN [4]: a representative node comparison-based method that generates histogram features based on pairwise node similarity scores.
  • GMN-match [16]: a representative node comparison-based method that alternatively propagates node information within and between graphs to update node information and implicitly establish node mapping.
  • GMN-emb [16]: a representative graph comparison-based method which generated graph representation using gate-weighted node representations.
  • GraphSim [5]: a representative node comparison-based method that reorders the similarity matrix by performing BFS traversal to capture the local affinity of matches.
  • GOTSim [9]: a representative node comparison-based method that leverages linear assignment solver to explicitly establish node mapping at each layer.
  • TaGSim [3]: a representative operation type-aware graph-level comparison method.
  • ERIC [35]: the state-of-the-art graph comparison-based method proposed to train models with node-graph similarity in an unsupervised manner, and decouple this module in the inference.
  • Greed [22]: a method handles GED computation by computing the Euclidean distance between graph representations.
  • GEDGNN [20]: the recent node comparison-based method that explicitly builds node mapping by optimizing the graph transport costs.
  • GraphEdX [13]: the state-of-the-art node comparison-based method that defines a family of distance measures for each operation types and reweighs the total cost with operation costs.
We generally use the official implementation and hyperparameters provided by the authors. The implementations of VJ and Beam are sourced from [29]. For models predicting the normalized GED value, we remove their last sigmoid layer to ensure a fair comparison.
C.3 Implementation Details
We apply a 5-layer Graph Isomorphism Network (GIN) [28] with jump connections as GNN encoder, whose hidden state and output are of 64 dimensionality. All MLPs in this work have adopted layer normalization and applied ReLU as the activation function. The outputs of these MLPs are of a dimensionality of 32. We used Adam as the optimizer, with a fixed learning rate of 0.001.
C.4 Experimental Protocol
All models are trained on a NVIDIA A6000 GPU with an AMD EPYC 7282 CPU. The batch size is set to 128 graph pairs for training and 2048 for validation and test. Due to the large number of possible graph pairs, each epoch consists of 100 iterations, except for synthetic graphs where pair counts are limited. Each model undergoes ten warm-up epochs and is evaluated on the validation set at the end of each epoch. Early stopping with a patience of 50 is applied, terminating training if the validation loss does not improve over 50 consecutive validation steps, to prevent overfitting while ensuring sufficient training.
C.5 Evaluation Metrics
Following previous work, we evaluate models based on prediction error and ranking ability. Prediction error measures the point-wise difference between predicted and ground-truth GED values. At the same time, ranking ability is assessed by simulating retrieval scenarios in graph databases, where training graphs serve as data graphs, test graphs serve as queries. During the testing phase, each model calculates the GED between the query graph and every graph in the database. In this context, the model should return the most relevant data graphs with the lowest GED for each query. We use Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) to assess GED prediction errors. Similar to MSE, RMSE is sensitive to outliers but presents errors in the original units for better interpretability, while MAE captures general errors.
For ranking performance, we use Spearman’s Rank Correlation Coefficient ($\rho$), Kendall’s Rank Correlation Coefficient ($\tau$), and Precision at $k$ ($p@k$). $\rho$ is computed as:
$\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)},$
where $d_{i}$ represents the difference in ranks for each data point between the two rankings, and $n$ is the total number of data points, the factor of 6 is included to normalize the sum of squared rank differences, ensuring that $\rho$ falls within the range $[-1,1]$ [25]. This metric measures the strength and direction of monotonic relationships. $\tau$ evaluates ordinal consistency and is computed as:
$\tau=\frac{C-D}{\binom{n}{2}},$
where $C$ is the number of pairs ranked in the same order in both rankings, $D$ is the number of pairs ranked in opposite orders in the two rankings, $\binom{n}{2}$ is the total number of possible pairs in a set of $n$ items. $p@k$ checks the relevance of the top $k$ predicted graphs against the ground truth GED ranking for each query. It calculates the intersection of the predicted top $k$ data graphs and the ground-truth top $k$ results, dividing the count of the intersection by $k$ for each query, and then averaging the results across all queries. To evaluate the ranking ability, we sorted all data graphs for each query based on both the ground truth and predicted GED values in ascending order, and then compared the resulting rankings.
Appendix D Results
D.1 Effectiveness
We provide additional analysis of the baseline models’ performance across datasets:
  • GraphEdX performs well on the node-labeled AIDS700nef dataset by leveraging XOR operations as node-level similarity. However, it underperforms on the unlabeled LINUX dataset and runs out of memory on IMDBMulti due to its design choice of padding all graphs to the same size, rather than within graph pairs like GEDGNN and Gen.
  • GMN-match fuses and propagates node representations across graphs in an iterative manner. While it performs competitively on AIDS700nef and LINUX, it suffers from error accumulation caused by its greedy fusion process.
  • GraphSim employs BFS-based matrix reordering to improve structural alignment. This design, however, introduces instability, as reflected by large standard deviations across runs.
Figure 7: Case Study of ranking results on AIDS700nef (labeled) and IMDBMulti (unlabeled).
D.2 Case Study
Figure 7 presents a case study comparing the top-ranked graphs retrieved by Gen and the ground truth GED rankings computed by A*. The example demonstrates that Gen closely approximates the true ranking, successfully identifying and ranking structurally similar graphs.
D.3 Performance on Larger Synthetic Graphs
Additional experimental in Figure 8 show that GMN-match produced higher outlier predictions. While SimGNN’s predictions generally align with the true distribution of GED, they are less concentrated than ours and exhibit more outlier predictions. We did not report results for another graph comparison method, Eric, because it proved particularly challenging to train due to gradient exploding issues. Despite our efforts to implement layer normalization at each layer of Eric, these attempts did not alleviate the problem. Aligned with other graph comparison models, Greed does not perform as well as its node comparison counterparts.
Figure 8: Performance of models on larger synthetic graphs across diverse node size groups.
D.4 Hyperparameter Sensitivity